You roll dice twice. Find the probability of the compound event "The first roll was even AND the second roll was larger than 4"
Answer by Boreal(15235)

Two ways to interpret this problem.
One die rolled twice: (1/2) probability even* (1/3) larger than 4 (5 and 6)=1/6 total probability.
----------------------
Two dice (because plural form of noun die).
First roll is even
That occurs when 2,4,6,8,10,12, and that is (1/2) as well but second part, greater than 4 is 30/36 or 5/6. That joint probability is (1/2)*(5/6)=(5/12)
I suspect the first solution is what is asked for.